OpenStudy (anonymous):

Solve the equation for y. 4x – 5y = 9

OpenStudy (arabpride):

4x – 5y = 9 -4x -4x 5y = -4x + 9 /5 /5 y = -4/5 x + 9/5

OpenStudy (anonymous):

Thank you!

OpenStudy (arabpride):

You're Welcome!! *^.^*

OpenStudy (anonymous):

y=4x/5 - 9/5....... thats what i got